OpenStudy (anonymous):

x+3+3x=7+4 step by step rv2 column proofs please?

OpenStudy (anonymous):

i don't understand the def. of sub. and stuff can someone please explain?

OpenStudy (john_es):

You should try to put the "x" in one side, and the numbers in other side. For example, \[x+3x=7+4-3\] In order to put the 3 in the right side, you have to deduct it from both sides.

OpenStudy (anonymous):

what are the stepsd called like commutative ...... & defintion of subtration

OpenStudy (john_es):

Well, I don't think I can say the steps in the same way they teach you, but I call the first pass, "substraction", because we substract a 3 from boths sides.

OpenStudy (john_es):

The you have to apply summation, substraction and conmutative properties (supposing they are steps), \[4x=11-3\Rightarrow 4x=8\Rightarrow x=2\] The last step, I think it should be called substitution.